OpenStudy (anonymous):

Write 0.42424242... as a ratio of two integers in lowest terms

OpenStudy (anonymous):

Could you show us what you've done so far please?

OpenStudy (anonymous):

If you haven't done anything, here's a Hint. Let x=0.42424242... Find 100x=?

OpenStudy (anonymous):

so 42? im just kind of lost

OpenStudy (anonymous):

What's \[100\times 0.4242424242...=?\]

OpenStudy (anonymous):

42.4242424242

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now to eliminate the repeaters, we subtract 100x with x. \[100x-x=99x\] So what would: \[99x=?\]

OpenStudy (anonymous):

41.9998 which would be 42?

OpenStudy (anonymous):

Uhmmm, it's just 42. It's not 41.9998. You shouldn't use a calculator if you are using one.

OpenStudy (anonymous):

Because both x and 100x have the same repeating units after the decimal place and that makes clear sense when you subtract both of them together, you would end up with a whole number.

OpenStudy (anonymous):

Do you understand that?

OpenStudy (anonymous):

OOOOO that makes a lot more sence now. Thanks so much

OpenStudy (anonymous):

sense*

OpenStudy (anonymous):

No worries, now you find what x equals which is what you're supposed to find but in whole numbers.\[99x=42\] what would: \[x=?\]

OpenStudy (anonymous):

@MrMcNuggets Are you there?

OpenStudy (anonymous):

ya just thinking, last statement somewhat confused me again lol, im terrible at math

OpenStudy (anonymous):

x=42/99?

OpenStudy (anonymous):

Oh okay. At the very start we let x=0.42424242....And the question asked us to make 0.42424242... into a ratio of two whole numbers.

OpenStudy (anonymous):

And yes, that's correct!

OpenStudy (anonymous):

Now the question wants the fraction in a ratio.

OpenStudy (anonymous):

ok! lol, i thought 42/99 was too easy to be the answer

OpenStudy (anonymous):

That will be very simple. Just take the two numbers: 42 and 99 and put them together with a colon in between the two numbers.

OpenStudy (anonymous):

And just to make sure you understand, 42/99 is not the answer.

OpenStudy (anonymous):

42:99 is the complete answer.

OpenStudy (anonymous):

42:99

OpenStudy (anonymous):

Yep. Fantastic. Well done!

OpenStudy (anonymous):

ok gotcha, thank you so much, i understand it better now

OpenStudy (anonymous):

take care

OpenStudy (anonymous):

No worries mate. Glad I could help. Cy.

OpenStudy (anonymous):

Cya*